# PLRD: Partially Linear Regression Discontinuity Inference

### v1

@@ 1 Introduction
The fractionWe are grateful to Evan Munro for sharing results on the behavior of existing methods for RDDs under generative adversarial simulation following Athey et al. (2024), and to P. Aronow for drawing our attention to their behavior on pure noise. This work was supported by the National Science Foundation under grant SES-2242876, the Office of Naval Research under grants N00014-17-1-2131 and N00014-19-1-2468, and gift from Amazon. of empirical NBER working papers using regression discontinuity designs (RDDs) increased from nearly zero in 1995 to more than 10% by 2015 (Currie, Kleven, and Zwiers, 2020, Figure 4B). Despite the ubiquity of RDDs in applied work, however, there are still open questions as to econometric best practices for inference in RDDs. In particular, although standard interpretation of inferential statements relies on confidence intervals achieving nominal coverage of true treatment effects, we find that widely used methods systematically fail to achieve nominal coverage.
rdrobust (b.c.) rdrobust (rob.) RDHonest plrd
@@
Lee (2008) 91.0% 0.036 95.4% 0.042 97.6% 0.052 98.3% 0.035
Cattaneo et al. (2015) 90.7% 5.353 95.1% 6.281 97.4% 6.980 97.3% 4.525
Matsudaira (2008) (read.) 93.8% 0.061 96.9% 0.069 98.4% 0.096 97.1% 0.060
Matsudaira (2008) (math) 88.6% 0.045 93.0% 0.054 97.8% 0.072 97.0% 0.051
Jacob and Lefgren (2004) (read.) 88.4% 0.067 95.0% 0.080 98.2% 0.104 96.5% 0.065
Jacob and Lefgren (2004) (math) 87.8% 0.078 93.4% 0.093 97.1% 0.120 95.0% 0.069
Oreopoulos (2006) 84.1% 0.196 94.6% 0.264 98.9% 0.316 96.1% 0.194
Ludwig and Miller (2007) 91.2% 4.058 95.7% 4.786 97.6% 5.238 97.4% 3.542
Meyersson (2014) 89.5% 6.362 93.3% 7.531 96.3% 8.099 97.3% 5.095
@@
Table 1: Empirical coverage (EC) and width of 95% confidence intervals from different methods, on simulated RDDs calibrated to nine applications. Following Athey et al. (2024) we generate our simulation distributions by training a Wasserstein Generative Adversarial Neural Network (WGAN) on the original data. We report results for rdrobust (Calonico et al., 2014) (both the “b.c.” method justified by asymptotics and the “rob” method with a finite-sample variance correction) and RDHonest (Armstrong and Kolesár, 2018). We also report results for our proposed plrd method. We highlight (in green) the entries whose coverage is not statistically significantly (i.e., at the 5% level using a one-sided binomial test) below nominal 95% coverage and (in blue) the entries with the shortest widths among those delivering valid coverage. All results are aggregated across 1000 simulation replications.
@@
We assess the performance of existing methods by evaluating them on simulations calibrated to nine high-profile applications of RDDs. We focus on approaches proposed by Calonico, Cattaneo, and Titiunik (2014) and Armstrong and Kolesár (2018), as implemented in the R functions rdrobust and RDHonest respectively. These methods are representative of current state of the art for inference in RDDs (Stommes, Aronow, and Sävje, 2023), and both offer automated end-to-end confidence interval constructions that can directly be applied to raw data. The method rdrobust by Calonico et al. (2014) is the dominant approach to regression-discontinuity inference in current econometric practice. [^1] This is the only general purpose regression discontinuity package listed in the “Econometrics Task View” for highly used R packages on CRAN (https://cran.r-project.org/web/views/Econometrics.html); and both the R (Calonico et al., 2015) and Stata (Calonico et al., 2017) versions of the function are widely used and cited. The method starts by running a local linear regression as recommended by Hahn, Todd, and der Klaauw (2001), and then applies a 2nd-order bias correction to mitigate curvature bias. The method rdrobust offers two options for building confidence intervals—a basic (b.c.) method which inherits the standard error estimator from the original local linear regression (and is justified by asymptotic arguments), and a robust (rob.) method that incorporates a finite-sample variance inflation to account for the bias correction—and we consider both. Meanwhile, the method RDHonest by Armstrong and Kolesár (2018) explicitly inflates the width of local linear regression confidence intervals to accommodate worst-case plausible curvature bias (rather than trying to correct for this bias). [^2] In principle, the method RDHonest asks the user to specify a worst-case bound on the curvature of the underlying conditional response functions up front. However, the method also has a default fall-back option where curvature bounds are derived using the rule-of-thumb of Armstrong and Kolesár (2020); in our experiments we use this default.
To assess these methods, we use generative adversarial neural networks to generate simulation instances that are representative of real-world applications, as recommended in Athey, Imbens, Metzger, and Munro (2024). [^3] The generative adversarial learning approach provides a simulator whose data-generating distribution closely mimics the moments and idiosyncracies of the original data (see Fig. 1 for an illustration). The generated simulator does not use an explicit parametric model and does not enforce any built-in smoothness assumptions, thus bringing it towards a neutral comparison ground for evaluating methods that were motivating under different assumptions. We summarize results for 95% confidence intervals built using the considered methods in Table 1; see Section 4 for details on the simulation design. A first observation is that rdrobust systematically undercovers the true regression discontinuity parameter. The basic (b.c.) intervals only achieve results consistent with nominal coverage in 1 out of 9 settings. The finite-sample adjusted intervals (rob.) are wider, but still only achieve nominal coverage in 6 out of the 9 settings. These qualitative findings do not appear to be specific to our simulation designs; for example, in Table 2 we observe similar coverage properties when applying the method to pure noise with no relationship between the outcome and the running variable.
rdrobust (b.c.) rdrobust (rob.) RDHonest plrd
Dataset EC width EC width EC width EC width
pure noise 88.7% 0.997 92.8% 1.181 94.9% 1.293 96.3% 0.847
[…]
@@
Meanwhile, the method RDHonest achieves valid coverage, but at the cost of substantially wider intervals: The median ratio between the length of these intervals and the basic bias-corrected (b.c.) ones varies between 1.27 to 1.60 in our simulations. Furthermore, these confidence intervals are often excessively conservative, thus suggesting that some of this additional width isn’t strictly needed. [^4] In several settings, RDHonest achieves 98% coverage or above—and this may come at a substantial cost in width. For example, in the classical setting with no bias adjustment, Gaussian 98% confidence intervals are 19% wider than the conventional 95% intervals. Our findings leave open a pragmatic question: Is it possible to design automated confidence intervals that reliably achieve nominal coverage in realistic settings, all while matching rdrobust in terms of interval length?
@@
The main contribution of this paper is a method, partially linear regression discontinuity inference (PLRD), that provides an affirmative answer to this question—at least in the context of the numerical settings we consider. As described further below, algorithmically, the PLRD estimator adapts the method of Armstrong and Kolesár (2018) to accommodate a partially linear treatment effect model; conceptually, however, our full approach fits within the econometric framework of Calonico et al. (2014). We implement our method in the package plrd for R (R Core Team, 2023), which provides end-to-end functionality for regression discontinuity inference without relying on the user to specify hard-to-choose tuning parameters. As seen in Tables 1 and 2, plrd achieves nominal and near-exact coverage throughout, and in every instance yields the shortest intervals among those with valid coverage.
@@ 2 The PLRD Estimator
    Under Assumption 1 there exists a finite constant $B$ such that the second derivatives $\mu^{\prime\prime}_{w}(x)$ are $B$-Lipschitz in a neighborhood of $c$; and it is possible to identify such a $B$ from data in the large-sample limit. Following Imbens and Kalyanaraman (2012) and Calonico et al. (2014), PLRD starts by estimating $B$ via a cubic regression.
@@
Our full procedure—including some extra algorithmic details such as cross fitting—is detailed as Algorithm 1. Our main formal result about the PLRD estimator is that it yields valid confidence intervals under our assumptions in the large-sample limit.
@@
[…]
@@
2:Randomly split the data into two near-equally sized folds, with index sets $I_{1}$ and $I_{2}$.
3:Use the observations in fold $I_{2}$ to estimate the Lipschitz constant $B$ by fitting cubic polynomials before and after the threshold as in (9), and denote this estimate as $\widehat{B}_{(2)}$.
4:Enforce a lower bound $\widehat{B}_{(2)}\leftarrow\max\{\widehat{B}_{(2)},\,\varepsilon\}$.
5:Estimate $\sigma^{2}$ by running the ordinary least-squares regression of $Y_{i}$ on the interaction of $X_{i}$ and $W_{i}$ for $i\in I_{2}$. Denote this estimate as $\widehat{\sigma}^{2}_{(2)}$ and record the residuals from this regression as:
@@
$(\widehat{\gamma}_{(1)},\widehat{t}_{(1)}):=\mathop{\rm{minimize}}_{\gamma,t}\left(\widehat{B}^{2}_{(2)}t^{2}+\widehat{\sigma}^{2}_{(2)}\sum_{i\in I_{1}}\gamma_{i}^{2}\right)\quad\text{subject to}\quad\sup_{\rho\in\mathcal{R}}\ \sum_{i\in I_{1}}\gamma_{i}\rho(X_{i}-c)\leq t.$ (3)
7:Repeat the steps 3 to 5 with the roles of $I_{1}$ and $I_{2}$ swapped.
8:Set $\widehat{\gamma}_{i}=\gamma_{(k),i}/2$ if $i\in I_{k}$ ($k=1,2$), and $\widehat{\gamma}_{i}=0$ if $|X_{i}-c|>\ell$. Compute $\widehat{\tau}_{\texttt{PLRD}}=\sum_{i=1}^{n}\widehat{\gamma}_{i}Y_{i}$, and $\widehat{s}^{2}(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}r_{i}^{2}$ where the $r_{i}$ are the residuals from the least squares regressions in step 4.
9:Report the following $100(1-\alpha)\%$ confidence interval for $\tau_{c}$:
@@
where $Z\sim\mathcal{N}(0,1)$, and $\widehat{b}:=(\widehat{B}_{(2)}\widehat{t}_{(1)}+\widehat{B}_{(1)}\widehat{t}_{(2)})/2$.
@@
Consider a sharp regression discontinuity design under Assumptions 1 and 2, where the running variable $X_{i}$ has a density strictly positive at the threshold $c$, $\sigma_{i}^{2}={\rm Var}(Y_{i}\mid X_{i})$ is bounded away from 00, i.e., $\sigma_{i}\geq\sigma_{\min}$ for some $\sigma_{\min}>0$, and $\mathbb{E}[|Y_{i}-\mu_{W_{i}}(X_{i})|^{q}\mid X_{i}=x]\leq C$ uniformly over all $x$, for some $q>2$ and $C\in(0,\infty)$. Suppose furthermore that we execute the initial cubic regression used to pick $B$ over a pilot bandwidth of length $\ell$ with
@@
and choose any $\varepsilon>0$ in Algorithm 1. Then, level-$(1-\alpha)$ PLRD intervals $\mathcal{I}_{\alpha}$ are asymptotically valid: For any $10<\alpha<$ 1,
@@
We note that our algorithm does in principle require two tuning parameters, namely the pilot bandwidth $\ell$ used to run the initial cubic regression to choose $B$ and the lower bound $\varepsilon$ for $\widehat{B}$. We note, however, that our algorithmic performance is largely insensitive to these choices (and simple defaults appear to provide good performance in practice); and similar objects are also required in Calonico et al. (2014) and Imbens and Kalyanaraman (2012). First, regarding $\ell$, this bandwidth can in practice be chosen simply to be “large”: We only need $\ell$ to decay faster than $n^{-1/12}$, whereas the mean-squred error optimal bandwidth for local linear regression in our setting (i.e., under Assumption 1) scales as $n^{-1/7}$ (Cheng, Fan, and Marron, 1997). In our implementation, we by default simply set $\ell$ to contain the full range of the data, and this is choice is what we use in all our experiments. [^6] This choice mirrors Imbens and Kalyanaraman (2012, p. 942, step 2) who seed their bandwidth-selection procedures by fitting global cubic polynomials to the full dataset; see also Calonico et al. (2014, S.2.6, step 0). As such, practitioners should be able to get reliable results out of our method unless the nature of the relationship between $X_{i}$ and $Y_{i}$ changes completely as we get far from the cutoff—and in settings like this it would be possible to pick $\ell$ by plotting the data. Meanwhile, $\varepsilon$ will have no effect in large samples as long as we choose it to be small enough that $\varepsilon<|\mu^{\prime\prime\prime}(c)|$. [^7] One interesting question is what happens at the critical point where $\mu^{\prime\prime\prime}(c)=0$; as, in this case, our constraint does necessarily bind and we fall back to using $\varepsilon$ as our curvature bound. The challenge is that, when $\mu^{\prime\prime\prime}(c)=0$, our cubic polynomial estimator $\widehat{B}$ (in step 3 of Algorithm 1) converges to 0, but the exact way it does so is left underspecified under Assumption 1. Our use of the lower bound $\varepsilon$ guarantees validity of our confidence intervals at the expense of some excess conservativeness when $\mu^{\prime\prime\prime}(c)=0$. We note that existing methods, including Calonico et al. (2014) and Imbens and Kalyanaraman (2012), also do not offer guidance on how to potentially benefit from “superefficiency” in similar critical cases; i.e., our approach is again in line with current best practices.

### v2

@@ 1 Introduction
The fractionWe are grateful to Evan Munro for sharing results on the behavior of existing methods for RDDs under generative adversarial simulation following Athey et al. (2024). We are also grateful to Tim Armstrong, P. Aronow, Max Farrell, Paul Goldsmith-Pinkham, Michal Kolesár, Christoph Rothe and Rocío Titiunik for insightful conversations and for comments on a previous version of this paper. This work was supported by the National Science Foundation under grant SES-2242876, the Office of Naval Research under grants N00014-17-1-2131 and N00014-19-1-2468, and a gift from Amazon. Software and replication files are available at https://github.com/ghoshadi/plrd. of empirical NBER working papers using regression discontinuity designs (RDDs) increased from nearly zero in 1995 to between 15% and 20% by 2015 (Currie et al., 2020, Figure 4B). [^1] Goldsmith-Pinkham (2024) maintains up-to-date numbers at https://paulgp.com/econlit-pipeline. Despite the ubiquity of RDDs in applied work, however, there are still open questions as to econometric best practices for inference in RDDs. In particular, we find that widely used methods systematically fail to achieve nominal coverage.
conventional rdrobust RDHonest plrd
@@
Lee (2008) 94.4% 0.037 94.8% 0.044 97.3% 0.056 98.1% 0.045
Cattaneo et al. (2015) 94.6% 5.003 95.0% 5.884 97.5% 6.796 96.1% 3.870
Ludwig and Miller (2007) 94.4% 5.954 94.9% 7.053 97.8% 7.808 96.8% 4.272
Meyersson (2014) 93.8% 6.301 94.5% 7.431 96.8% 7.869 96.3% 4.618
Matsudaira (2008) (read.) 86.4% 0.050 92.2% 0.059 97.9% 0.103 97.7% 0.059
Matsudaira (2008) (math) 92.0% 0.046 93.4% 0.054 98.3% 0.094 97.6% 0.052
Jacob and Lefgren (2004) (read.) 94.2% 0.068 94.3% 0.082 97.9% 0.114 97.5% 0.076
Jacob and Lefgren (2004) (math) 94.4% 0.059 95.3% 0.071 98.2% 0.124 97.9% 0.061
Oreopoulos (2006) 93.5% 0.189 94.6% 0.255 98.8% 0.303 96.2% 0.124
@@
Table 1: Empirical coverage (EC) and (average) width of 95% confidence intervals from different methods, on simulated RDDs calibrated to nine applications. Following Athey et al. (2024) we generate our simulation distributions by training a Wasserstein Generative Adversarial Neural Network (WGAN) on the original data. We report results for the conventional local linear regression with bandwidth chosen as recommended in Imbens and Kalyanaraman (2012); robust bias-corrected (rdrobust) confidence intervals of Calonico et al. (2014); and bias-aware (RDHonest) confidence intervals as proposed in Armstrong and Kolesár (2018). We also report results for our proposed plrd method. We highlight (in green) the entries whose coverage is not statistically significantly (i.e., at the 5% level using a one-sided binomial test) below nominal 95% coverage and (in blue) the entries with the shortest widths among those delivering valid coverage. All results are aggregated across 10,000 simulation replications.
@@
We assess the performance of three commonly used methods by evaluating them on simulations calibrated to nine high-profile applications of RDDs. First, we consider conventional local linear regression (Hahn, Todd, and der Klaauw, 2001) with bandwidth chosen as in Imbens and Kalyanaraman (2012). We also consider two more recent proposals by Calonico, Cattaneo, and Titiunik (2014) and Armstrong and Kolesár (2018), as implemented in the R functions rdrobust and RDHonest respectively. These methods are representative of the current state of the art for inference in RDDs (Stommes, Aronow, and Sävje, 2023), and both offer automated end-to-end confidence interval constructions that can directly be applied to raw data. The method rdrobust by Calonico et al. (2014) is a dominant approach to regression-discontinuity inference in current econometric practice. [^2] This is the only general purpose regression discontinuity package listed in the “Econometrics Task View” for highly used R packages on CRAN (https://cran.r-project.org/web/views/Econometrics.html); and both the R (Calonico et al., 2015) and Stata (Calonico et al., 2017) versions of the function are widely used and the original paper (Calonico et al., 2014) cited is over 4500 times as of now. The method starts by running a local linear regression as recommended by Hahn, Todd, and der Klaauw (2001), and then applies a 2nd-order bias correction to mitigate curvature bias. [^3] The method rdrobust offers two options for building confidence intervals: A basic bias-corrected method which inherits the standard error estimator from the original local linear regression (and is justified by asymptotic arguments), and a robust method that incorporates a finite-sample variance inflation to account for the bias correction. The latter is the recommended one in the paper and this is the one we consider here. Meanwhile, the method RDHonest by Armstrong and Kolesár (2018) explicitly inflates the width of local linear regression confidence intervals to accommodate worst-case plausible curvature bias (rather than trying to correct for this bias). [^4] In principle, the method RDHonest asks the user to specify a worst-case bound on the curvature of the underlying conditional response functions up front. However, the method also has a default fall-back option where curvature bounds are derived using the rule-of-thumb of Armstrong and Kolesár (2020); in our experiments we use this default.
To ensure that the evaluations of these methods are credible, we follow Athey, Imbens, Metzger, and Munro (2024) and use generative adversarial neural networks (Goodfellow et al., 2014) to generate simulation instances that are representative of a set of seven seminal regression discontinuity studies. [^5] Two of the studies have two outcomes, so we have a total of nine applications. The generative adversarial learning approach yields simulators whose data-generating distribution closely mimics the moments and idiosyncrasies of the original datasets; see Figure 1 for an illustration. The generated simulators do not use explicit parametric models and do not enforce any built-in smoothness assumptions, thus bringing us towards a neutral comparison ground for evaluating methods that were motivated under different assumptions.
This semi-synthetic simulation setup is an important part of our argument. It comes with two qualifications. First, it relies on the availability of datasets where regression discontinuity methods are applicable. We apply it to nine such data sets, but we encourage researchers to apply it to additional ones. Second, it relies on our generative adversarial nets yielding realistic data generating processes given the provided datasets. Well-known results in non-parametric inference imply that, fixing a sample size and given any dataset where standard regression-discontinuity estimators work well, one can construct another adversarial dataset that is statistically indistinguishable from the first but where the same estimators fail to get coverage (Low, 1997). Our approach to evaluation implicitly relies on an assumption that the real datasets we are calibrated to aren’t adversarial in this sense. Furthermore, our estimation process involves some choices of tuning parameters for which different researchers may make different choices; in our experiments we seek to mitigate concerns of researcher bias by using a single set of tuning parameters across all datasets.
@@
We summarize results for 95% confidence intervals in Table 1; see Section 4 for details on the simulation design. A first observation is that the conventional local linear regression intervals systematically undercover the true parameter. The method rdrobust achieves better—although not perfect—coverage, with nominal rates in 4 out of the 9 settings. However, it does so at the expense of wider confidence intervals than conventional local linear regression (18% to 36% wider in our applications). RDHonest in turn achieves reliable coverage throughout all designs, but at the cost of still-wider intervals: The median ratio between the width of these intervals and the rdrobust ones varies between 1.06 to 1.76 in our simulations, and between 1.25 to 2.11 relative to the conventional ones. Furthermore, these confidence intervals are often excessively conservative, thus suggesting that some of this additional width is not strictly needed. [^6] In several settings, RDHonest achieves 98% coverage or above—and this may come at a substantial cost in width. For example, in the classical setting with no bias adjustment, Gaussian 98% confidence intervals are 18.7% wider than the conventional 95% intervals. These qualitative findings do not appear to be specific to our simulation designs; for example, in Table 3 we observe similar coverage properties when applying the method to pure noise with no relationship between the outcome and the running variable.
These semi-synthetic simulation findings leave open a pragmatic question: Is it possible to design automated confidence intervals that reliably achieve nominal coverage in realistic settings, while remaining competitive with the conventional intervals in terms of interval width? We emphasize here the “realistic” qualifier. Even if it is not feasible to achieve uniformly better properties, it may be that most real-world applications are sufficiently smooth to allow for nominal coverage and narrower confidence intervals than the existing methods.
@@
The main contribution of this paper is a method, partially linear regression discontinuity inference (PLRD), that provides an affirmative answer to this question—at least in the context of the numerical settings we consider. As described further below, algorithmically, the PLRD estimator adapts the method of Armstrong and Kolesár (2018) to accommodate a partially linear treatment effect model; conceptually, however, our full approach fits within the econometric framework of Calonico et al. (2014). We implement our method in the package plrd for R (R Core Team, 2023), which provides end-to-end functionality for regression discontinuity inference without relying on the user to specify hard-to-choose tuning parameters. As illustrated in Tables 1 and 3, our proposed plrd method achieves nominal and near-exact coverage throughout, and in nearly every instance yields the shortest intervals among those with valid coverage.
@@ 2 The PLRD Estimator
    Under Assumption 1 there exists a finite constant $B$ such that the second derivatives $\mu^{\prime\prime}_{w}(x)$ are $B$-Lipschitz in a neighborhood of $c$; and it is possible to identify such a $B$ from data in the large-sample limit. Following Imbens and Kalyanaraman (2012) and Calonico et al. (2014), PLRD starts by estimating $B$ via a cubic regression. [^7] We use sample splitting Algorithm 1 to justify technical arguments underlying formal results given in Section 5. It is possible, however, that an alternative argument could be used to justify similar results without sample splitting; see, e.g., Armstrong and Kolesár (2020, Appendix E) for results of this type.
@@
Our full procedure—including some extra algorithmic details such as cross fitting—is detailed as Algorithm 1. As a pragmatic safeguard against failures of Assumption 2, we start our procedure with a hypothesis test, and if we reject this assumption with high confidence then we proceed with a variant of our method that only relies on Assumption 1; see Appendix A for further discussion.
@@
Tuning Parameters: Window size $\ell$ for estimating the smoothness parameter $B$ (default is set as the whole range), a lower bound $\varepsilon$ for $\widehat{B}$ (default is $\varepsilon={\operatorname{SD}[Y]}\,/\,100$), and a confidence level $1-\alpha^{\prime}$ for curvature detection (default is set as $99.9\%$).
@@
2:Perform an ANOVA test between cubic polynomials fitted with or without different curvatures at significance level $1-\alpha^{\prime}$ to detect whether different curvatures should be used.
3:Randomly split the data into two near-equally sized folds, with index sets $I_{1}$ and $I_{2}$.
4:Estimate $\sigma^{2}$ by running the ordinary least-squares regression of $Y_{i}$ on the interaction of $X_{i}$ and $W_{i}$ for $i\in I_{2}$. Denote this estimate as $\widehat{\sigma}^{2}_{(2)}$ and record the residuals from this regression as:
@@
5:if the test for different curvatures in Step 2 is not rejected then
6:  Use the observations in fold $I_{2}$ to estimate the Lipschitz constant $B$ by fitting cubic
7:   polynomials with the same curvature as in (10), and denote this estimate as $\widehat{B}_{(2)}$.
8:  Enforce a lower bound $\widehat{B}_{(2)}\leftarrow\max\{\widehat{B}_{(2)},\,\varepsilon\}$.
9:  Let $\mathcal{R}:=\{f:f(0)=f^{\prime}(0)=f^{\prime\prime}(0)=0,f^{\prime\prime}\text{ is }1\text{-Lipschitz}\}$, and solve the following optimization problem via quadratic programming (see Appendix D for details):
@@
10:  Repeat Steps 6 to 9 with the roles of $I_{1}$ and $I_{2}$ swapped.
11:else
12:  Use the observations in fold $I_{2}$ to estimate the Lipschitz constant $B$ by fitting cubic
13:   polynomials with different curvatures (cf. Appendix A), and denote this estimate as $\widehat{B}_{(2)}$.
14:  Enforce a lower bound $\widehat{B}_{(2)}\leftarrow\max\{\widehat{B}_{(2)},\,\varepsilon\}$.
15:  Let $\mathcal{R}:=\{f:f(0)=f^{\prime}(0)=f^{\prime\prime}(0)=0,f^{\prime\prime}\text{ is }1\text{-Lipschitz}\}$, and solve the following optimization problem via quadratic programming (see Appendix D for details):
[…]
16:  Repeat Steps 12 to 15 with the roles of $I_{1}$ and $I_{2}$ swapped.
17:end if
18:Set $\widehat{\gamma}_{i}=\gamma_{(k),i}/2$ if $i\in I_{k}$ ($k=1,2$), and $\widehat{\gamma}_{i}=0$ if $|X_{i}-c|>\ell$. Compute $\widehat{\tau}_{\texttt{PLRD}}=\sum_{i=1}^{n}\widehat{\gamma}_{i}Y_{i}$, and $\widehat{s}^{2}(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}r_{i}^{2}$ where the $r_{i}$ are the residuals from the least squares regressions in (2).
19:Report the following $100(1-\alpha)\%$ confidence interval for $\tau_{c}$:
@@
where $Z\sim\mathcal{N}(0,1)$, and $\widehat{b}:=\big(\widehat{B}_{(2)}\widehat{t}_{(1)}+\widehat{B}_{(1)}\widehat{t}_{(2)}\big)/2$.
Our main formal result about the PLRD estimator is that it yields valid confidence intervals under our assumptions in the large-sample limit. [^9] In our formal results, we assume a continuous running variable with support near the cutoff. In the case of discrete running variables, steps 2 and 3 of our procedure remain valid, i.e., given an appropriate $B$ they yield partial identification intervals for $\tau_{c}$ with a guarantee of the type (7) (Imbens and Wager, 2019; Kolesár and Rothe, 2018). However, when $X_{i}$ has discrete support, step 1 is no longer guaranteed to consistently recover $B$ under Assumption 1 alone; and so further assumptions would be required to justify data-driven detection of $B$.
@@
Consider a sharp regression discontinuity design under Assumptions 1 and 2, where the running variable $X_{i}$ has a strictly positive density at the threshold $c$, $\sigma_{i}^{2}={\rm Var}(Y_{i}\mid X_{i})$ is bounded away from 0, i.e., $\sigma_{i}\geq\sigma_{\min}$ for some $\sigma_{\min}>0$, and $\mathbb{E}[|Y_{i}-\mu_{W_{i}}(X_{i})|^{q}\mid X_{i}=x]\leq C$ uniformly over all $x$, for some $q>2$ and $C\in(0,\infty)$. Suppose furthermore that we execute the initial cubic regression used to pick $B$ over a pilot bandwidth of width $\ell$ with
@@
[…]
@@
We note that our algorithm does in principle require two tuning parameters, namely the pilot bandwidth $\ell$ used to run the initial cubic regression to choose $B$ and the lower bound $\varepsilon$ for $\widehat{B}$. We note, however, that our algorithmic performance is largely insensitive to these choices (and simple defaults appear to provide good performance in practice); and similar objects are also required in Calonico et al. (2014) and Imbens and Kalyanaraman (2012). First, regarding $\ell$, this bandwidth can in practice be chosen simply to be “large”: We only need $\ell$ to decay faster than $n^{-1/12}$, whereas the mean-squared error optimal bandwidth for local linear regression in our setting (i.e., under Assumption 1) scales as $n^{-1/7}$ (Cheng, Fan, and Marron, 1997). In our implementation, we by default simply set $\ell$ to contain the full range of the data, and this choice is what we use in all our experiments. [^10] This choice mirrors Imbens and Kalyanaraman (2012, p. 942, step 2) who seed their bandwidth-selection procedures by fitting global cubic polynomials to the full dataset; see also Calonico et al. (2014, S.2.6, step 0). As such, practitioners should be able to get reliable results out of our method unless the nature of the relationship between $X_{i}$ and $Y_{i}$ changes completely as we get far from the cutoff—and in settings like this it would be possible to pick $\ell$ by plotting the data. Meanwhile, $\varepsilon$ will have no effect in large samples as long as we choose it to be small enough that $\varepsilon<|\mu^{\prime\prime\prime}(c)|$. [^11] One interesting question is what happens at the critical point where $\mu^{\prime\prime\prime}(c)=0$; as, in this case, our constraint does necessarily bind and we fall back to using $\varepsilon$ as our curvature bound. The challenge is that, when $\mu^{\prime\prime\prime}(c)=0$, our cubic polynomial estimator $\widehat{B}$ (in step 3 of Algorithm 1) converges to 0, but the exact way it does so is left underspecified under Assumption 1. Our use of the lower bound $\varepsilon$ guarantees validity of our confidence intervals at the expense of some excess conservativeness when $\mu^{\prime\prime\prime}(c)=0$. We note that existing methods, including Calonico et al. (2014) and Imbens and Kalyanaraman (2012), also do not offer guidance on how to potentially benefit from “superefficiency” in similar critical cases; i.e., our approach is again in line with current best practices.